OpenStudy (anonymous):

Determine on which intervals the function is increasing and decreasing. f(x) = -(x - 9)2 + 8

ganeshie8 (ganeshie8):

sketching it may help

OpenStudy (anonymous):

Ok so first, Alyy, is it f(x) = -(x - 9)2 + 8 or f(x) = -(x - 9)^2 + 8 Remember the ^ sign changes everything like I told you before

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So does it have a ^ or no? f(x) = -(x - 9)2 + 8 or \[f(x) = -(x - 9)^{2} + 8\]

OpenStudy (anonymous):

The second equation needs the ^2 to denote it is squared.

OpenStudy (anonymous):

- infinity, 9?

OpenStudy (anonymous):

increasing?

OpenStudy (anonymous):

Ok it is asking anyone who clicks that to login. Do not ever give your login information to anyone ok Alyy?

OpenStudy (anonymous):

ohh ok oops

OpenStudy (anonymous):

i wont

OpenStudy (anonymous):

K so, the graph will look like this, 

OpenStudy (anonymous):

We need to know the x-intercepts and y-intercept to know where it is increasing and decreasing from. Do you know how to find the x and y-intercepts using the equation?

OpenStudy (anonymous):

yea if the graph was numbered

OpenStudy (anonymous):

Sorry, not the y-intercept, but its maximum value